OpenStudy (kalijah13):

What expression shows the relationship between the value of any term and n, its position in the sequence for the given sequence?

OpenStudy (kalijah13):

10, 5, 0, -5, -10, ... A) -5n B) -5n + 15 C) -3n + 15 D) 5n

OpenStudy (kalijah13):

I WILL ALSO AWARD A MEDAL AND POSSIBLY A FAN

OpenStudy (welshfella):

this is an arithmetic sequence in which the common difference is -5

OpenStudy (welshfella):

the general formula for each term in this sequence is a1 +d(n -1) where a1 = first term , n = sequence number and d = common difference here a1 = 10 and d = -5

OpenStudy (kalijah13):

Thank you so much

OpenStudy (welshfella):

sorry i was wrong there the correct choice is -5n + 15

OpenStudy (kalijah13):

okay

OpenStudy (welshfella):

a1 + d(n -1) = a1 + dn - d when a1 = 10 and d = -5 we have 10 -5n - -5 = -5n + 15

OpenStudy (kalijah13):

okay